# Interpretable Responsibility Sharing as a Heuristic for Task and Motion Planning

### v1

@@ 1 Introduction
Figure 1: This study emphasizes the common human factor in both environment construction and task specifications present in household tasks (e.g., serving, cleaning, and caring), and how it influences Task and Motion Planning (TAMP). We highlight the role of human bias, represented through the use of auxiliary objects such as trays in kitchens, which are often created by humans for convenience. Our research proposes a systematic method to utilize these biases in an interpretable manner, aiming to enhance the efficiency and effectiveness of TAMP in human-designed environments.
@@
Domestic robots are expected to handle household tasks defined by humans to improve their standards of living. These tasks include serving, cleaning, and caring in environments also created by humans. Due to this common human factor in sequential decision-making and manipulation problems, incorporating human bias into robots’ decision-making processes can enhance their overall performance [1, 2] (Figure 1). Task and Motion Planning (TAMP) can effectively address these household tasks. Current TAMP research  [3, 4, 5] often focuses on formulating these high-level tasks at a logical level, generating task plans using either uninformed or informed search, and finding feasible motions to execute these plans in a sequential or interleaved fashion. However, these studies often overlook human bias.
Our perspective on human bias in this study is that when humans construct their environment, they also create auxiliary objects that might be beneficial for certain human-defined tasks. For example, a tray often exists in a kitchen environment and is frequently used as the initial object for a household serving task. Therefore, we propose utilizing these objects when necessary, as they reflect human bias in these situations. We are not concerned with any other potential influences or interventions possibly made by humans. Our goal is to develop a systematic method to leverage this bias, achieving a more efficient and effective TAMP formulation.
Figure 2: We propose a novel heuristic, Interpretable Responsibility Sharing (IRS), for Task and Motion Planning (TAMP). We begin by curating a dataset consisting of first-order logic definitions of initial states and goals, with labels derived from counterfactual scenarios indicating whether the use of auxiliary objects would benefit the agent. The Optimized Rule Synthesis (ORS) component then generates a set of rules that define the appropriate conditions for the agent to use these objects. If the initial state and goal do not violate these rules, the agent adopts responsibility sharing by dividing the original problem into sub-problems dependent on the auxiliary objects. IRS enhances the effectiveness and interpretability of existing TAMP formulations by incorporating human bias through the ORS-guided decision mechanism.
Several methods have been developed to tackle sequential decision-making and manipulation problems using TAMP  [6, 7, 8, 9, 10, 11, 12]. While recent approaches using heuristics perform well  [13, 14, 15, 16], the heuristic functions are often constrained to specific task domains, limiting their generalizability to other robotic tasks and their application in real-world household scenarios. Uninformed task planners  [17, 18], which generate plans by performing a tree search through a knowledge base, are more generalizable due to their domain-agnostic nature. However, they struggle to find plans in a reasonable time frame because of the combinatorial nature of the search problem, where the search space increases exponentially, reducing their applicability  [16]. Additionally, they aim to find feasible plans with the minimum number of high-level actions, which may lead to sub-optimal plans.
Moreover, both approaches fail to leverage the human factor. While this is not a concern in uninformed search, heuristic-defined, informed search-based TAMP formulations often adopt an oracle approach. This oracle informs the planner about the cost of an action, typically based on displacement, which may not accurately reflect human preferences. Recognizing the importance of the human factor, researchers have recently explored using annotated data or structured reward mechanisms to train deep policies with sequential planning objectives  [19, 20, 21, 22]. These methods show promising results in both performance and adaptability to unseen environments, making them a viable alternative for real-world robotic applications. Furthermore, they implicitly incorporate human bias through curated datasets or tailored reward mechanisms. However, the application of deep policy methods in task and motion planning faces interpretability issues, making embodied agents without a transparent decision-making process potentially unsafe for use in households  [23, 24].
@@
Therefore, effective and interpretable task and motion planning require meeting two main criteria: leveraging the human bias present in sequential decision-making and manipulation problems and enhancing interpretability in the decision-making process.
@@
To meet these criteria, we propose a novel planning heuristic, Interpretable Responsibility Sharing (IRS), which leverages human bias while providing an interpretable perspective on how an embodied agent uses this heuristic in its decision-making process (Figure 2). To address this inductive bias—referring to the assumptions or patterns derived from human-designed environments and tasks that influence the agent’s behavior or decision-making process—we introduce a new concept called Responsibility Sharing (RS).
@@
In household tasks, auxiliary objects such as trays often exist in the environment for serving items. The robot can manipulate the target objects either directly with its end-effector or indirectly using these auxiliary objects. We refer to this indirect relationship as Responsibility Sharing. By delegating some responsibilities to auxiliary objects (e.g., placing the objects on a tray, which the agent then carries), the original problem is divided into sub-problems (move the objects to the tray, then move the tray to the final destination). This approach simplifies and enhances the effectiveness of the combined planning problem by leveraging inductive bias.
It should be noted that this concept differs from tool usage, as using these objects is not mandatory to complete the task. For example, one may choose to use a tray or not while serving. This non-mandatory nature of these objects presents challenges for existing TAMP formulations. In uninformed approaches, using these objects increases the depth of the search, while in existing informed approaches, the task-dependent nature limits the range of objects that can be effectively utilized.
To address this non-mandatory nature, we need an adaptive yet interpretable decision-making mechanism to determine when to use these auxiliary objects across multiple tasks. We propose a novel approach called Optimized Rule Synthesis (ORS), which generates rules for agents to decide whether to utilize non-mandatory objects in the environment, based on state information defined under first-order logic.
ORS operates through two fundamental components: Representative Rule-Based Learner (RRL) [25] and Correlation and Order-Aware Rule Learning (CARL) [26]. These components are vital for the functionality of ORS: RRL is crucial for creating and learning rules based on system data, identifying patterns and relationships between actions and object states, and establishing the foundational rules that guide initial decision-making processes. Enhancing ORS, CARL introduces sensitivity to the sequence of actions and the relationships between different object states and actions, ensuring that the rules maintain logical coherence while adapting effectively to dynamic environments.
By integrating these critical components, ORS constructs a robust system capable of making informed predictions and justifying the use of auxiliary objects in various scenarios, thereby enhancing the adaptive capabilities of interpretable domestic robots.
ORS is trained on a counterfactual dataset created through Counterfactual Plan Generation (CPG), ensuring that auxiliary objects are responsible for performance changes. This training enables ORS to identify when and how to use these objects in an interpretable manner, enhancing transparency and decision-making in real-world applications. By combining ORS and RS, IRS guides the TAMP formulation, improving effectiveness and efficiency in an interpretable manner, particularly in domains where data is scarce (e.g., robotics). Overall, this work introduces the following contributions:
@@
    Interpretable Responsibility Sharing (IRS), a heuristic for task and motion planning, aiming to improve the effectiveness and interpretability of the agent by leveraging the inherent human bias present in sequential decision-making and manipulation problems.
@@
    Optimized Rule Synthesis (ORS) is a novel rule-based decision-making mechanism that determines when an agent should share responsibilities with auxiliary objects based on environmental settings across multiple tasks.
@@
We evaluate our IRS through three TAMP tasks suitable for domestic robots: serving, pouring, and handover. In the serving task, objects are carried to their final destinations, while in the pouring task, a source is distributed to objects at different locations using auxiliary objects like trays and pitchers. The handover task offers a unique perspective on the boundaries of responsibility sharing, as a stationary robot acts as an auxiliary object, demonstrating the applicability of IRS in multi-robotic settings. A human experiment is conducted to observe the presence of Responsibility Sharing (RS) within human decision-making mechanisms. Our ablation studies provide valuable insights into decision-making performance and interpretability. These findings, validated through both qualitative and quantitative evaluations, underscore the effectiveness and interpretability of our approach.
@@ 2.1 Task and Motion Planning (TAMP)
Combined TAMP involves jointly solving high-level symbolic action planning and low-level motion planning  [3, 9, 27]. TAMP settings are defined by a combination of symbolic actions, states, and physical constraints, with tasks represented by either symbolic or physical goals. To identify feasible high-level action sequences for these goals, various TAMP planners have been developed. These planners have evolved from linear, flow-like approaches  [28, 29] to more complex, interleaved planners  [4, 5, 30, 31] that integrate motion planning to varying extents, ensuring the feasibility of symbolic plans. However, TAMP inherently faces scalability challenges due to large action and state spaces, limiting the horizon of proposed plans and their adaptability to complex scenarios  [32, 33]. Common strategies to address this include heuristic-guided search  [16, 34, 35] and breaking down the original problem into sub-problems  [36, 37, 38]. Our approach aims to divide the original problem into sub-problems using the proposed concept of responsibility sharing, if it is inferred to improve the agent’s performance.
@@ 2.2 Rule-Based and Advanced Knowledge Graph Reasoning Methods
Rule-Based: Traditional rule-based models such as decision trees, rule lists, and rule sets, while being transparent, often face limitations in terms of scalability and generalization [39, 40, 41]. These models rely heavily on heuristic methods or prolonged itemset mining, which may not always yield the most effective solutions and are time-intensive on larger datasets. To enhance these, ensemble models like Random Forests and Gradient Boosted Decision Trees have been employed, which typically offer better performance by integrating multiple sub-models [42]. However, the complexity of these ensemble methods often detracts from their interpretability, which is crucial for certain applications such as robotics [43].
Gradient-Based Discrete Model Training: Addressing the challenges of training rule-based models, gradient-based methods for discrete model training like the Straight-Through Estimator (STE) and Gradient Grafting have emerged [44, 45]. These methods attempt to optimize binary or quantized neural networks, enhancing both network compression and acceleration. The integration of gradient data from both discrete and continuous models aims to improve the efficacy of training processes without sacrificing the benefits of rule-based logic [46].
Inductive Logic Programming (ILP) extends rule-based modeling by utilizing induction to derive logical rules from observed instances and background knowledge, aiming to generate understandable and verifiable rules [47]. However, similar to traditional rule-based approaches, ILP struggles with scalability and data noise. Path-based reasoning, another facet of symbolic logic-based methods, deduces general laws from specific observations but often lacks generalization capabilities, despite producing precise and interpretable results [48, 49].
Advanced Knowledge Graph Reasoning Methods:  The embedding of knowledge graphs represents a significant shift towards computational efficiency in reasoning methodologies. Techniques like TransE, DistMult, and RotateE encode entities and relations into vectors within a defined space, facilitating fast and effective reasoning processes [50, 51, 52]. However, these embedding methods generally suffer from a lack of interpretability, presenting challenges for applications that require transparent decision-making processes. In contrast, neural-symbolic methods blend the computational prowess of neural networks with the logical rigor of symbolic reasoning [53]. This hybrid approach not only enhances model performance and efficiency but also retains a degree of interpretability through the use of symbolic constraints on embeddings.
The evolution from traditional rule-based models to advanced neural-symbolic and embedding-based techniques illustrates a broader trend in knowledge graph reasoning towards integrating computational efficiency with logical robustness. Each methodology compensates for the limitations of its predecessors, striving to balance performance with interpretability. For instance, while ILP and traditional rule-based models provide a solid foundation in logical reasoning, their scalability issues are addressed by ensemble methods.Further, the computational limitations observed in ensemble and traditional methods are mitigated by embedding and neural-symbolic techniques, which also attempt to reintroduce interpretability lost in purely computational approaches. Our approach with IRS aims to leverage the strengths of both symbolic reasoning and machine learning methods, focusing on creating models that are not only effective and efficient but also transparent and adaptable to a wide range of tasks.
@@ 2.3 Interpretability in Robotics
Among these, fuzzy rule-based methods utilize fuzzy logic to manage uncertainty and vagueness in robotic decision-making [57, 58, 59]. These methods generate fuzzy rules either from expert knowledge or through data-driven learning. However, the interpretability of fuzzy rules may diminish as the number of rules and the complexity of their membership functions increase [60]. Furthermore, although fuzzy logic adeptly handles partial truths, its nuanced nature can obscure the decision-making process, especially when compared to more structured, binary approaches [61].
In contrast, Graph Neural Networks (GNNs) offer a promising alternative for managing complex robotic manipulation tasks by utilizing relational data, which potentially enhances the interpretability of decisions [62, 63, 64]. GNNs model task-relevant entities and their interrelations, supporting generalizable and somewhat interpretable robotic tasks. Nevertheless, the effectiveness of GNNs is often contingent upon the availability of substantial data sets, and their computational demands can be prohibitive, especially for real-time applications in robotics.
Our method, termed Interpretable Responsibility Sharing (IRS), overcomes these challenges by focusing on a human-centric design that capitalizes on inherent biases and common practices in human environments. IRS employs a rule-based framework that is computationally more efficient than the graph-based calculations required by GNNs, facilitating its real-time application in robotic systems. By decomposing complex tasks into simpler sub-tasks and employing Optimized Rule Synthesis (ORS) for rule generation, IRS ensures that each decision is both transparent and effective. This structured simplicity and focus on human-centric design not only enhance interpretability but also make IRS particularly well-suited for task and motion planning in domestic settings.
@@ 3 Background
Interpretable Responsibility Sharing (IRS) serves as a heuristic for Task and Motion Planning. In this study, we used Logic Geometric Programming (LGP) and Multi-Bound Tree Search (MBTS) to solve TAMP problems  [4, 65]. During our dataset construction, we created counterfactual scenarios and used Individual Treatment Effect to determine the better scenario  [66]. Representative Rule-Based Learner (RRL)  [25] and Correlation and Order-Aware Rule Learning (CARL)  [26] are the main components of ORS, providing interpretable responsibility sharing. This section will provide a brief background on these concepts
@@ 3.1 Logic-Geometric Programming
Logic Geometric Programming (LGP) serves as the foundational framework for solving TAMP problems. We will now outline the optimization strategy of LGP, as formulated in Equation 1  [4]. Let the configuration space $X\subset\mathbb{R}^{n}\times SE(3)^{m}$ represent $m$ rigid objects and $n$ articulated joints, starting from the initial condition $x_{0}$. The objective of LGP is to optimize a sequence of symbolic actions $a_{1:K}$, states $s_{1:K}$, and the associated continuous trajectory $x(t)$, where $t\in\mathbb{R}$ maps to $X$, in order to achieve a symbolic goal $g$. Positions, velocities, and accelerations are denoted by $\bar{x}=(x,\dot{x},\ddot{x})$. The domain of symbolic states $s\in S$ and actions $a\in A(s)$ is discrete and finite, with state transitions $s_{k-1}\to s_{k}$ defined by a first-order logic language, similar to PDDL.
[…]
@@
For any given sequence of actions $a_{1:K}$ and states $s_{1:K}$, the functions $h(\cdot)$, $g(\cdot)$, and $c(\cdot)$ are continuous and have segments that are piecewise differentiable. Optimizing the trajectory is an undertaking of a nonlinear program (NLP). At the heart of LGP lies the integration of discrete symbolic search (i.e., the pursuit of a sequence of symbolic states that culminate in the goal) with simultaneous nonlinear optimization. This combined strategy is tasked with computing trajectories that adhere to the defined constraints and determining the feasibility of actualizing a sequence of symbolic actions in a continuous manner.

### v2

@@ 1 Introduction
Figure 1: The influence of Human-Centric Environmental Bias on TAMP.
@@
This study emphasizes the common human factor in both environment construction and task specifications present in household tasks (e.g., serving, cleaning, and caring). We highlight the role of Human-Centric Environmental Bias, represented through the use of auxiliary objects such as trays in kitchens, which are often created by humans for convenience. Our research proposes a systematic method to utilize these biases in an interpretable manner, aiming to enhance the efficiency and effectiveness of TAMP in human-designed environments.
@@
Domestic robots are expected to handle household tasks defined by humans to improve their standards of living. These tasks include serving, cleaning, and caring in environments also created by humans. Due to this common human factor in sequential decision-making and manipulation problems, incorporating human-centric environmental bias into robots’ decision-making processes can enhance their overall performance [46, 28] (Figure 1).
In this study, we introduce the term Human-Centric Environmental Bias to describe the inherent structure embedded in domestic settings. When humans construct their environments, they introduce auxiliary objects (e.g., trays, pitchers) specifically designed to facilitate recurring tasks. Unlike approaches that leverage human bias through imitation learning from demonstrations, our perspective focuses on the inductive bias present in the environment itself. We propose that the mere presence of these human-designed objects encodes a preferred strategy (e.g., a tray implies a grouping strategy for serving). Our goal is to develop a systematic method to interpret and leverage these environmental cues, thereby achieving a more efficient TAMP formulation without requiring explicit human demonstrations.
Task and Motion Planning (TAMP) can effectively address these household tasks [11, 32, 25, 17, 22, 52, 6]. Current TAMP research  [19, 60, 21] often focuses on formulating these high-level tasks at a logical level, generating task plans using either uninformed or informed search, and finding feasible motions to execute these plans in a sequential or interleaved fashion.
However, these studies often overlook human-centric environmental bias due to the limitations of existing approaches. Uninformed task planners [45, 7] are domain-agnostic but suffer from exponential search spaces, limiting their applicability. Conversely, heuristic-based approaches [54, 16, 67, 4] often rely on oracle cost functions based on low-level physical metrics (e.g., displacement or energy), which fail to capture human preferences or the strategic utility of auxiliary objects. While recent deep learning methods [61, 65, 10, 64] show promise in adaptability by implicitly learning these biases from data, they lack transparency, making the resulting agents potentially unsafe for domestic deployment [37, 36].
@@
Therefore, effective and interpretable task and motion planning require meeting two main criteria: leveraging the human-centric environmental bias present in sequential decision-making and manipulation problems and enhancing interpretability in the decision-making process.
Figure 2: The Interpretable Responsibility Sharing (IRS) Framework. The pipeline consists of three core phases: (1) Data Generation, where Counterfactual Plan Generation (CPG) creates a labeled dataset based on physical effort; (2) Rule Synthesis, where Optimized Rule Synthesis (ORS) learns interpretable logical conditions for auxiliary object usage; and (3) Heuristic Planning, where IRS utilizes these rules to guide the TAMP solver by dynamically decomposing tasks into sub-problems.
@@
To meet these criteria, we propose a novel planning heuristic, Interpretable Responsibility Sharing (IRS), which leverages human-centric environmental bias while providing an interpretable perspective on how an embodied agent uses this heuristic in its decision-making process (Figure 2). Crucially, IRS acts as a meta-heuristic that is orthogonal to existing TAMP solvers; it guides the search by dividing the problem into sub-problems independent of the underlying exploration strategy.
Unlike traditional heuristics that compute low-level physical costs (e.g., displacement) online to guide the search, IRS utilizes these metrics strictly as offline supervisory signals during the learning phase. By compiling physical costs into symbolic rules, our approach shifts the computational complexity of physical evaluation from inference time to the training phase, aligning with hierarchical sub-goal decomposition methods [14].
The core pipeline of our framework operates in three stages to bridge environmental bias and planning efficiency:
  1. 1.
    Data Generation (CPG): We employ Counterfactual Plan Generation to create a dataset that labels whether using an auxiliary object improves performance for a given task configuration.
  2. 2.
    Rule Synthesis (ORS): Using this dataset, Optimized Rule Synthesis learns interpretable logical rules that predict when auxiliary objects should be used.
  3. 3.
    Heuristic Planning (IRS): During inference, IRS uses these rules to guide the TAMP solver. If the rules trigger Responsibility Sharing (RS)—the delegation of transport tasks to auxiliary objects—the planner decomposes the problem into efficient sub-goals; otherwise, it defaults to standard search.
@@
While the term Responsibility Sharing may evoke concepts from multi-agent systems [23], we extend it here to the relationship between an agent and auxiliary objects. By delegating some responsibilities to auxiliary objects (e.g., placing the objects on a tray, which the agent then carries), the original problem is divided into sub-problems. This approach simplifies and enhances the effectiveness of the combined planning problem by leveraging inductive bias. It should be noted that this concept differs from tool usage, as using these objects is not mandatory to complete the task.
This study specifically aims to answer how embodied agents can systematically interpret and leverage these inductive biases, such as the strategic use of auxiliary objects, to enhance planning efficiency. Our primary focus is to develop an interpretable heuristic that identifies when sharing responsibility is beneficial by compiling physical intuition into high-level logical rules. To isolate and validate this decision-making logic, we limit our experimental scope to fully observable settings and assume sufficient physical capacity for auxiliary tools. This abstraction allows us to evaluate the effectiveness of the proposed heuristic in capturing human-centric environmental bias without the confounding complexity of partial observability or dynamic bin-packing.
Overall, this work introduces the following contributions:
@@
    Interpretable Responsibility Sharing (IRS), a heuristic for task and motion planning, aiming to improve the effectiveness and interpretability of the agent by leveraging the inherent human-centric environmental bias.
@@
    Optimized Rule Synthesis (ORS), a novel rule synthesis framework that integrates feature-based and relation-aware generators through a bounded greedy selection process. ORS introduces a Balance Score objective to filter rules based on their informational sufficiency, ensuring the synthesized policy captures the necessary physical context to ground robotic decision-making.
@@
We evaluate our IRS through three TAMP tasks suitable for domestic robots: serving, pouring, and handover. In the serving task, objects are carried to their final destinations, while in the pouring task, a source is distributed to objects at different locations using auxiliary objects like trays and pitchers. The handover task offers a unique perspective on the boundaries of responsibility sharing, as a stationary robot acts as an auxiliary object, demonstrating the applicability of IRS in multi-robotic settings. A human experiment is conducted to observe the presence of Responsibility Sharing (RS) within human decision-making mechanisms. Our ablation studies provide valuable insights into decision-making performance and interpretability.
@@ 2.1 Task and Motion Planning (TAMP)
Combined TAMP involves jointly solving high-level symbolic action planning and low-level motion planning  [19, 17, 12]. TAMP settings are defined by a combination of symbolic actions, states, and physical constraints, with tasks represented by either symbolic or physical goals. To identify feasible high-level action sequences for these goals, various TAMP planners have been developed.
Early approaches often relied on linear, flow-like architectures [24, 44]. Modern solvers have evolved into complex, interleaved planners that generally fall into two distinct categories. Sampling-based methods, such as PDDLStream [21] and FFRob [20], integrate symbolic planning with geometric reasoning by treating geometric constraints as black-box samplers; they generate candidate values for continuous parameters during the search and verify feasibility through external motion planning queries. In contrast, optimization-based strategies, most notably Logic-Geometric Programming (LGP) [60], formulate the TAMP problem as a constrained mathematical program, optimizing symbolic choices and continuous trajectories simultaneously to satisfy physical constraints.
While these methods ensure physical consistency, TAMP inherently faces scalability challenges due to large action and state spaces, limiting the horizon of proposed plans and their adaptability to complex scenarios  [41, 13]. Common strategies to address this include heuristic-guided search  [4, 15, 53] and breaking down the original problem into sub-problems  [14, 59, 27].
@@ 2.2 Rule-Based and Advanced Knowledge Graph Reasoning Methods
Rule-Based and Gradient Methods: Traditional rule-based models such as decision trees and rule sets, while transparent, often face limitations in scalability. To enhance performance, ensemble methods like Random Forests have been employed, but their complexity often detracts from the interpretability required for robotics [26]. While Inductive Logic Programming (ILP) [40] offers verifiable logical rules, it struggles with scalability on noisy data. To address these limitations, gradient-based methods for discrete model training, such as the Straight-Through Estimator (STE) and Gradient Grafting [8, 62], have emerged. These techniques optimize discrete logical structures using continuous gradient information, bridging the gap between symbolic rigor and neural scalability.
Knowledge Graph Reasoning: In the domain of knowledge graphs, embedding techniques like TransE and RotateE [3, 55] facilitate fast reasoning but suffer from opacity. In contrast, neural-symbolic methods blend the computational prowess of neural networks with the logical rigor of symbolic reasoning [48], retaining interpretability through symbolic constraints. Our approach with IRS leverages this synergy, combining gradient-based rule learning (RRL) with order-aware graph reasoning (CARL) to ensure both effectiveness and transparency.
@@ 2.3 Interpretability in Robotics
Among these, fuzzy rule-based methods utilize fuzzy logic to manage uncertainty and vagueness in robotic decision-making [51, 1, 39]. Unlike fuzzy systems, where interpretability is often inversely proportional to the number of rules due to the ambiguity of overlapping membership functions, our approach leverages First-Order Logic (FOL) to define crisp, binary boundaries. While fuzzy interpretability suffers when the rule base becomes dense, interpretability in Task and Motion Planning benefits from descriptiveness. Effective robotic planning requires rules that possess sufficient conditions to ground the action in physical reality (e.g., specifying locations or object counts) [56]. Thus, distinct from fuzzy logic where less is more prevents cognitive overload, we posit that for robotic agency, sufficient is more is required to ensure causal validity [33].
Alternatively, Graph Neural Networks (GNNs) offer a promising alternative for managing complex robotic manipulation tasks by utilizing relational data, which potentially enhances the interpretability of decisions [35, 34, 31]. GNNs model task-relevant entities and their interrelations, supporting generalizable and somewhat interpretable robotic tasks. Nevertheless, the effectiveness of GNNs is often contingent upon the availability of substantial data sets, and their computational demands can be prohibitive, especially for real-time applications in robotics.
@@ 3 Background
Interpretable Responsibility Sharing (IRS) serves as a heuristic for Task and Motion Planning. In this study, we used Logic Geometric Programming (LGP) and Multi-Bound Tree Search (MBTS) to solve TAMP problems  [60, 58]. During our dataset construction, we created counterfactual scenarios and used Individual Treatment Effect to determine the better scenario  [43]. Representative Rule-Based Learner (RRL)  [63] and Correlation and Order-Aware Rule Learning (CARL)  [29] are the main components of ORS, providing interpretable responsibility sharing.
@@ 3.1 Logic-Geometric Programming
Logic Geometric Programming (LGP) serves as the foundational framework for solving TAMP problems. Let the configuration space $X\subset\mathbb{R}^{n}\times SE(3)^{m}$ represent $m$ rigid objects and $n$ articulated joints, starting from the initial condition $x_{0}$. The objective of LGP is to optimize a sequence of symbolic actions $a_{1:K}$, states $s_{1:K}$, and the associated continuous trajectory $x(t)$, where $t\in\mathbb{R}$ maps to $X$, in order to achieve a symbolic goal $g$. Positions, velocities, and accelerations are denoted by $\bar{x}=(x,\dot{x},\ddot{x})$. The domain of symbolic states $s\in S$ and actions $a\in A(s)$ is discrete and finite, with state transitions $s_{k-1}\to s_{k}$ defined by a first-order logic language.
$\hskip 28.45274pt\min_{x,s_{1:K},a_{1:K},K}\hskip 5.69054pt\int_{0}^{KT}c(\bar{x}(t),s_{k}(t))\,dt$ (1)
$\hskip 51.21495pt\text{s.t.}\hskip 2.84526pt\quad\bar{x}(0)=x_{0},$
@@
At the heart of LGP lies the integration of discrete symbolic search (i.e., finding a plan skeleton) with simultaneous nonlinear optimization. This combined strategy computes trajectories that adhere to constraints and determines the feasibility of symbolic actions.